OpenStudy (anonymous):

Find X if B is the midpoint of AC, AB= 3x+4 and BC= 12x-13

OpenStudy (anonymous):

well, if AC is the full length, then AB and BC must both be half length each (therefore, they are equal) so we can write this as: AB = BC 3x+4 = 12x-13 3x = 12x - 17 -9x = -17 x = 17/9 i dont have a calculator on me so i'll let you punch that in and there's your answer. :)

OpenStudy (anonymous):

thanks i've been stuck on this question for awhile, so i was just working on other assignments

OpenStudy (anonymous):

no worries, does it make more sense to you now? let me know if there's anything you didn't understand and i'll gladly go over it for you.

OpenStudy (anonymous):

how long are you usually on here?

OpenStudy (anonymous):

usually a couple of hours each weekday. basically whenever i'm at school in a class i can use my laptop without getting busted.

OpenStudy (anonymous):

cool